# Uncertainty-driven Embedding Convolution

### v2

@@ Abstract
Text embeddings are essential components in modern NLP pipelines. While numerous embedding models have been proposed, their performance varies across domains. This variability motivates the use of ensemble techniques to combine complementary strengths. However, most existing ensemble methods operate on deterministic embeddings and fail to account for model-specific uncertainty, limiting their robustness and reliability in downstream applications. To address these limitations, we propose Uncertainty-driven Embedding Convolution (UEC). UEC first transforms deterministic embeddings into probabilistic ones in a post-hoc manner. It then computes adaptive ensemble weights based on embedding uncertainty, grounded in a Bayes-optimal solution under a surrogate loss. Additionally, UEC employs an uncertainty-aware similarity function that directly incorporates uncertainty into the similarity scoring, providing a theoretically grounded and efficient surrogate to distributional distances. Extensive experiments on diverse benchmarks demonstrate that UEC consistently improves both performance and robustness by leveraging principled uncertainty modeling. [^1] Code is available at: https://github.com/MLAI-Yonsei/UEC.
@@ 1 Introduction
Embeddings are core building blocks in modern NLP, capturing semantic meaning for words, sentences, and documents. They support tasks like similarity (Gao et al., 2021), retrieval (MacAvaney et al., 2019), QA (Devlin et al., 2019), and classification (Cer et al., 2018). Numerous embedding models (Devlin et al., 2019; Liu et al., 2019; Reimers & Gurevych, 2019) have emerged with diverse architectures and training objectives, but their performance varies across tasks and domains. No single model excels universally; instead, models offer complementary strengths. This motivates combining multiple embeddings to leverage their diverse capabilities.
@@
Figure 1: Comparison of deterministic embedding ensemble and UEC. Deterministic averaging (left) uniformly averages embeddings without considering their reliability, often leading to suboptimal decisions. In this example, both candidate embeddings contribute equally, resulting in an incorrect retrieval. In contrast, the proposed UEC (right) adjusts weights based on the uncertainty, assigning higher importance to the more reliable candidate and successfully retrieving the correct answer.
@@
Embedding models can be ensembled at two levels: parameter and representation. At the parameter level, techniques like model merging combine model weights (Wortsman et al., 2022; Ilharco et al., 2022; Yang et al., 2024). However, this approach is often limited by strict architectural constraints. On the other hand, combining embeddings at the representation level, i.e., ensembling the output embeddings themselves, is a more practical and widely applicable strategy. However, most existing ensemble approaches rely on deterministic aggregation methods such as uniform averaging, which fail to consider the reliability or uncertainty of individual embeddings (Khan et al., 2023a; Fang et al., 2019). Without modeling uncertainty, these methods treat all embeddings as equally reliable, which can lead to suboptimal or unstable performance, especially when some models are poorly calibrated or mismatched to the target task. This is evident in Figure 1, where averaging conflicting ‘animal’ and ‘car’ interpretations for ‘jaguar’ leads directly to a retrieval failure.
@@
To address these limitations, we propose Uncertainty-driven Embedding Convolution (UEC), a framework for combining embeddings in a principled, uncertainty-aware manner. UEC consists of three key components. First, it converts pre-trained deterministic embeddings into probabilistic embeddings in a post-hoc fashion, allowing each embedding to represent both its mean and uncertainty. Second, it computes adaptive ensemble weights based on estimated uncertainty, down-weighting less reliable embeddings. This weighting strategy is grounded in a Bayes-optimal solution under a surrogate loss. Third, UEC introduces an uncertainty-aware similarity function that incorporates both distance and variance into the similarity score, offering a theoretically grounded and efficient surrogate to distributional distances. Consequently, empirical results on diverse benchmarks confirm that UEC reliably improves performance and robustness by leveraging principled uncertainty modeling.
@@
  • We introduce the UEC framework, which adaptively combines multiple embeddings by weighting their contributions based on estimated uncertainties, a method theoretically grounded as a Bayes-optimal solution under surrogate loss minimization.
@@ 2.1 Embedding Ensemble
Embedding ensemble methods aim to improve performance and robustness by combining multiple pre-trained models. Shuang et al. (2019) proposed CDWE, which handles polysemy by generating multiple word prototypes through deconvolution. Fang et al. (2019) enhanced knowledge-aware embeddings by re-weighting knowledge-graph edges. Liu et al. (2025) fused BERT (Devlin, 2018) variants using cross-attention guided by pseudo-labels. Sahlgren (2021) distilled multiple encoders into one by matching their averaged outputs. Khan et al. (2023b) explored fusion strategies for domain-specific BERT models, yielding minor improvements. Despite these advances, most approaches rely on deterministic embeddings, limiting their ability to model uncertainty. In contrast, our method ensembles probabilistic embeddings, capturing both distributional information and model diversity through principled uncertainty estimation.
@@ 2.2 Probabilistic Embedding
Probabilistic embeddings represent inputs as distributions rather than points, thereby modeling uncertainty in representation learning. Early approaches such as Gaussian (Vilnis & McCallum, 2014), Gaussian mixture (Chen et al., 2015), and Bayesian embeddings (Barkan, 2017) captured word-level distributional semantics. Sen2Pro (Shen et al., 2023) extended this idea to sentences using pre-trained models, but it requires fine-tuning, lacks post-hoc conversion, and does not support scalable similarity. Subsequent works explored multimodal and vision–language tasks: Chun et al. (2021); Chun (2023) improved cross-modal retrieval and image–text alignment, yet remained modality-specific; ProbVLM (Upadhyay et al., 2023) and BayesVLM (Baumann et al., 2024) incorporated uncertainty into frozen vision–language models through adapters or post-hoc strategies, but with limited generality. Other studies investigated hedging against ambiguous inputs (Oh et al., 2018), compositional multimodal retrieval (Neculai et al., 2022), uncertainty-aware multimodal pre-training (Ji et al., 2023), and scalable probabilistic embeddings with Gaussian process latent variable models (Venkataramanan et al., 2025). Our work provides a Gaussian convolution formulation that not only generates probabilistic embeddings in a post-hoc manner but also supports theoretically grounded and lightweight similarity estimation for downstream tasks.
@@ 3 Uncertainty-driven Embedding Convolution
We propose Uncertainty-driven Embedding Convolution (UEC), a principled framework for combining multiple embedding models by modeling predictive uncertainty. UEC consists of three stages (Figure 2): (1) post-hoc conversion of deterministic embeddings into probabilistic ones via Laplace approximation (LA) (Section 3.1), (2) Gaussian convolution with uncertainty-aware weights (Section 3.2), and (3) uncertainty-aware similarity scoring (Section 3.3). This enables robust and adaptive ensemble without manual tuning.
Figure 2: Overview of the UEC framework: UEC first transforms deterministic embeddings from multiple encoder models into probabilistic representations using Laplace approximation. These probabilistic embeddings are then adaptively combined by computing uncertainty-driven ensemble coefficients based on per-dimension variances. Finally, similarity is measured using a uncertainty-aware metric that accounts for both the mean and uncertainty of the ensembled embedding.
@@ 3.1 Post-hoc Probabilistic Embedding Model
We first transform deterministic embeddings into probabilistic embeddings using LA (MacKay, 1992; Ritter et al., 2018). This allows each embedding to express both mean and uncertainty, which we later exploit for adaptive model combination and similarity estimation.
@@
LA approximates the posterior distribution of model parameters $\mathbf{w}$ as a Gaussian centered at the maximum a posteriori (MAP) estimate. We denote the training dataset as $\mathcal{D}=\{(\mathbf{x},y)\}$, where $\mathbf{x}$ and $y$ are the input and corresponding output, respectively. The posterior is given by Bayes’ theorem:
$p(\mathbf{W}|\mathcal{D})=\frac{p(\mathcal{D}|\mathbf{W})p(\mathbf{W})}{p(\mathcal{D})},$
where $p(\mathcal{D}|\mathbf{W})$ is the likelihood, $p(\mathbf{W})$ is the prior, and $p(\mathcal{D})$ is the evidence. Since this posterior is generally intractable, LA approximates $p(\mathbf{W}|\mathcal{D})$ with a Gaussian centered at $\widehat{\mathbf{W}}$ (the MAP) and precision matrix given by the Hessian $\mathbf{H}_{\widehat{\mathbf{W}}}$:
[…]
$\mathbf{H}_{\widehat{\mathbf{W}}}$ $=\nabla^{2}[-\log p(\mathcal{D}|\mathbf{W})-\log p(\mathbf{W})]\Big|_{\mathbf{W}=\widehat{\mathbf{W}}}$.
This approximates posterior $p(\mathbf{W}|\mathcal{D})\approx\mathcal{N}(\widehat{\mathbf{W}},\mathbf{H}^{-1}_{\widehat{\mathbf{W}}})$.
@@
Since computing the full Hessian is often impractical for large models, recent work shows that applying diagonal LA to only the last layer yields sufficiently accurate uncertainty estimates with lower computational overhead (Daxberger et al., 2021). Following this line of work, we apply LA with a diagonal covariance approximation only to the final layer of the embedding model. Let $f(\mathbf{x})$ be an embedding model with $L$ layers, and let $\mathbf{W}^{(L)}$ and $\mathbf{h}^{(L-1)}$ denote the parameters and output of the penultimate layer, respectively. In the deterministic setting, the embedding is computed as $\mathbf{z}=\mathbf{W}^{(L)\top}\mathbf{h}^{(L-1)}$.
[…]
$\mathbf{z}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}(\mathbf{x}),\bm{\Sigma}(\mathbf{x}))=\mathcal{N}(\widehat{\mathbf{W}}^{(L)\top}\mathbf{h}^{(L-1)},\mathbf{h}^{(L-1)\top}\mathbf{H}^{-1}_{\widehat{\mathbf{W}}^{(L)}}\mathbf{h}^{(L-1)}).$
3.2 Bayes-optimal Weights for Uncertainty-driven Embedding Convolution
@@
Suppose we have $K$ independent embedding models, each producing a probabilistic embedding $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$. To integrate these into a single representation, we perform a Gaussian convolution in the embedding space. Formally, we define the convolutional embedding $\mathbf{z}(\mathbf{x})$ as the result of a weighted combination of independent Gaussian variables:
@@
Bayes-optimal Coefficients
To determine the optimal coefficients $\bm{\pi}_{k}(\mathbf{x})$ for Gaussian convolution, we define an uncertainty-aware surrogate loss that reflects how closely the embeddings of positive pairs are expected to align. We then derive a convex optimization problem that identifies the coefficients minimizing this surrogate loss.
Text embeddings are often trained with contrastive losses, such as InfoNCE (Oord et al., 2018). For $\ell_{2}$-normalized features, the squared Euclidean distance relates directly to cosine similarity: $\|\mathbf{u}-\mathbf{v}\|^{2}=2(1-\mathbf{u}^{\top}\mathbf{v})$ (Wang & Isola, 2020). These losses also encourage alignment of positive pairs and uniformity on the unit hypersphere. Leveraging this connection, we adopt the squared loss as a principled surrogate for contrastive objectives.
Let $(\mathbf{x},\mathbf{x}^{\prime})$ be a positive pair. For each model $k$ we observe $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$ and $\mathbf{z}_{k}(\mathbf{x^{\prime}})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x^{\prime}}),\bm{\Sigma}_{k}(\mathbf{x^{\prime}}))$. We define the uncertainty-aware surrogate loss, a multi-task objective that aggregates model-specific embedding errors weighted by uncertainty-aware task coefficients, as:
@@
This surrogate loss can be decomposed into fidelity and epistemic uncertainty. While $\bm{\mu}_{k}(\mathbf{x})$ and $\bm{\Sigma}_{k}(\mathbf{x})$ are directly computable from the input $\mathbf{x}$, the corresponding terms $\bm{\mu}_{k}(\mathbf{x}^{\prime})$ and $\bm{\Sigma}_{k}(\mathbf{x}^{\prime})$ require evaluating the embedding model at the reference point $\mathbf{x}^{\prime}$. However, in our retrieval setting—aligned with common practice—document embeddings must be precomputed and stored in advance, independent of the incoming query $\mathbf{x}^{\prime}$. This practical constraint makes it infeasible to rely on any query-dependent terms during document embedding construction. As such, we approximate the surrogate loss by retaining only the components that depend on $\mathbf{x}$, leading to the following approximation:
$\min_{\bm{\pi}}\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\left(\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\right)\quad\text{s.t.}\;\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=\mathbf{1},\;\bm{\pi}_{k}(\mathbf{x})\geq 0.$ (3)
This is a convex optimization problem over the probability simplex, where each model is penalized based on both its uncertainty and the magnitude of its embedding. Its unique closed-form solution is the inverse-variance weighting:
$\bm{\pi}_{k}^{\star}(\mathbf{x})\approx\frac{1/\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))}{\sum_{j=1}^{K}1/\operatorname{tr}(\bm{\Sigma}_{j}(\mathbf{x}))}.$ (4)
Eq. 4 generalized inverse-weighting scheme accounts for both epistemic uncertainty (via $\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))$ and representation fidelity (via $\|\bm{\mu}_{k}(\mathbf{x})\|^{2}$), allowing the ensemble to favor embeddings that are simultaneously confident and well-aligned with the input. This optimal coefficient $\bm{\pi}^{*}_{k}(\mathbf{x})$ minimizes the uncertainty-aware surrogate loss in Eq. 3, maintaining consistency with the original objective. Full derivation appears in Appendix A.1.

### v3

@@ Abstract
Text embeddings are essential components in modern NLP pipelines. Although numerous embedding models have been proposed, no single model consistently dominates across domains and tasks. This variability motivates the use of ensemble techniques to combine complementary strengths. However, most existing ensemble methods operate on deterministic embeddings and fail to account for model-specific uncertainty, limiting their robustness and reliability in downstream applications. To address these limitations, we propose Uncertainty-driven Embedding Convolution (UEC). UEC first transforms deterministic embeddings into probabilistic ones in a post-hoc manner. It then computes adaptive ensemble coefficients based on embedding uncertainty, derived from a principled surrogate-loss formulation. Additionally, UEC employs an uncertainty-aware similarity function that directly incorporates uncertainty into the similarity scoring, providing a theoretically grounded and efficient surrogate to distributional distances. Extensive experiments on diverse benchmarks demonstrate that UEC consistently improves both performance and robustness by leveraging principled uncertainty modeling. [^1] Code is available at: https://github.com/MLAI-Yonsei/UEC.
@@ 1 Introduction
Embeddings are core building blocks in modern NLP, capturing semantic meaning for words, sentences, and documents. They support tasks like similarity (Gao et al., 2021), retrieval (MacAvaney et al., 2019), QA (Devlin et al., 2019), and classification (Cer et al., 2018). Numerous embedding models (Devlin et al., 2019; Liu et al., 2019; Reimers and Gurevych, 2019) have emerged with diverse architectures and training objectives, but their performance varies across tasks and domains. No single model excels universally; instead, models offer complementary strengths. This motivates combining multiple embeddings to leverage their diverse capabilities.
@@
Figure 1: Comparison of deterministic embedding ensemble and UEC. Deterministic ensemble (left) uniformly averages embeddings without considering their reliability, often leading to suboptimal decisions. In this example, both candidate embeddings contribute equally, resulting in an incorrect retrieval. In contrast, the proposed UEC (right) adjusts weights based on the uncertainty, assigning higher importance to the more reliable candidate and successfully retrieving the correct answer.
@@
Embedding models can be ensembled at two levels: parameter and representation. At the parameter level, techniques like model merging combine model weights (Wortsman et al., 2022; Ilharco et al., 2022; Yang et al., 2024). However, this approach is often limited by strict architectural constraints. On the other hand, combining embeddings at the representation level, i.e., ensembling the output embeddings themselves, is a more practical and widely applicable strategy. However, most existing ensemble approaches rely on deterministic aggregation methods such as uniform averaging, which fail to consider the reliability or uncertainty of individual embeddings (Khan et al., 2023; Fang et al., 2019). Without modeling uncertainty, these methods treat all embeddings as equally reliable, which can lead to suboptimal or unstable performance, especially when some models are poorly calibrated or mismatched to the target task. This is evident in Figure 1, where averaging conflicting ‘animal’ and ‘car’ interpretations for ‘jaguar’ leads directly to a retrieval failure.
@@
To address these limitations, we propose Uncertainty-driven Embedding Convolution (UEC), a framework for combining embeddings in a principled, uncertainty-aware manner. UEC consists of three key components. First, it converts pre-trained deterministic embeddings into probabilistic embeddings in a post-hoc fashion, allowing each embedding to represent both its mean and uncertainty. Second, it computes adaptive ensemble weights based on estimated uncertainty, down-weighting less reliable embeddings. This weighting strategy is grounded in a principled, uncertainty-aware surrogate-loss formulation. Third, UEC introduces an uncertainty-aware similarity function that incorporates both distance and variance into the similarity score, offering a theoretically grounded and efficient surrogate to distributional distances. Consequently, empirical results on diverse benchmarks confirm that UEC reliably improves performance and robustness by leveraging principled uncertainty modeling.
@@
  • We introduce the UEC framework, which adaptively and data-dependently combines multiple embeddings by weighting their contributions based on estimated, query-specific uncertainties, following a principled surrogate-loss formulation.
@@ 2.1 Embedding Ensemble
Embedding ensemble methods aim to improve performance and robustness by combining multiple pre-trained models. Shuang et al. (2019) proposed CDWE, which handles polysemy by generating multiple word prototypes through deconvolution. Fang et al. (2019) enhanced knowledge-aware embeddings by re-weighting knowledge-graph edges. Liu et al. (2025) fused BERT (Devlin, 2018) variants using cross-attention guided by pseudo-labels. Sahlgren (2021) distilled multiple encoders into one by matching their averaged outputs. Khan et al. (2023) explored fusion strategies for domain-specific BERT models, yielding minor improvements. Despite these advances, most approaches rely on deterministic embeddings, limiting their ability to model uncertainty. In contrast, our method ensembles probabilistic embeddings, capturing both distributional information and model diversity through principled uncertainty estimation.
@@ 2.2 Probabilistic Embedding
Probabilistic embeddings represent inputs as distributions rather than points, thereby modeling uncertainty in representation learning. Early approaches such as Gaussian (Vilnis and McCallum, 2014), Gaussian mixture (Chen et al., 2015), and Bayesian embeddings (Barkan, 2017) captured word-level distributional semantics. Sen2Pro (Shen et al., 2023) extended this idea to sentences using pre-trained models, but it requires fine-tuning, lacks post-hoc conversion, and does not support scalable similarity. Subsequent works explored multimodal and vision–language tasks: Chun et al. (2021); Chun (2023) improved cross-modal retrieval and image–text alignment, yet remained modality-specific; ProbVLM (Upadhyay et al., 2023) and BayesVLM (Baumann et al., 2024) incorporated uncertainty into frozen vision–language models through adapters or post-hoc strategies, but with limited generality. Other studies investigated hedging against ambiguous inputs (Oh et al., 2018), compositional multimodal retrieval (Neculai et al., 2022), uncertainty-aware multimodal pre-training (Ji et al., 2023), and scalable probabilistic embeddings with Gaussian process latent variable models (Venkataramanan et al., 2025). Our work provides a Gaussian convolution formulation that not only generates probabilistic embeddings in a post-hoc manner but also supports theoretically grounded and lightweight similarity estimation for downstream tasks.
@@ 3 Uncertainty-driven Embedding Convolution
We propose Uncertainty-driven Embedding Convolution (UEC), a principled framework for combining multiple embedding models by modeling predictive uncertainty.
UEC performs three simple steps (Figure 2): (1) converts each deterministic embedding model into a probabilistic one that outputs a Gaussian embedding (Section 3.1), (2) combines these probabilistic embeddings so that more confident models contribute more (Section 3.2), and (3) computes an uncertainty-aware similarity score (Section 3.3). This enables a robust and adaptive ensemble without manual tuning.
Figure 2: Overview of the UEC framework: UEC first transforms deterministic embeddings from multiple encoder models into probabilistic representations using Laplace approximation. These probabilistic embeddings are then adaptively combined by computing uncertainty-driven ensemble coefficients based on per-dimension variances. Finally, similarity is measured using an uncertainty-aware metric that accounts for both the mean and uncertainty of the ensembled embedding.
@@ 3.1 Post-hoc Probabilistic Embedding Model
To convert a deterministic embedding model into a probabilistic one, we estimate a Gaussian posterior over its last-layer parameters using the Laplace Approximation (LA) (MacKay, 1992; Ritter et al., 2018). Given training data $\mathcal{D}$ and last-layer weights $\mathbf{W}^{(L)}$, LA constructs a second-order Taylor expansion of the negative log-posterior $\log p(\mathbf{W}^{(L)}|\mathcal{D})$ around the maximum a posteriori (MAP) estimate $\widehat{\mathbf{W}}^{(L)}$:
$\log p(\mathbf{W}^{(L)}|\mathcal{D})\approx-\log p(\widehat{\mathbf{W}}^{(L)}|\mathcal{D})+\frac{1}{2}(\mathbf{W}^{(L)}-\widehat{\mathbf{W}}^{(L)})^{\top}\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}(\mathbf{W}^{(L)}-\widehat{\mathbf{W}}^{(L)}),$
where $\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}$ is the Hessian of the negative log-posterior evaluated at the MAP. This MAP solution corresponds to the final-layer parameters of the pre-trained embedding model. Prior work shows that final-layer LA is effective and efficient (Daxberger et al., 2021; Hobbhahn et al., 2022), yielding a Gaussian weight posterior without retraining. The full derivation of LA is provided in Appendix A.1.
@@
Building on the LA posterior derived above, we convert the pre-trained embedding model into a probabilistic one in a post-hoc manner. Let $f(\mathbf{x})$ be an embedding model with last-layer parameters $\mathbf{W}^{(L)}$ and penultimate representation $\mathbf{h}^{(L-1)}(\mathbf{x})$. In the deterministic setting, the embedding is $\mathbf{z}(\mathbf{x})=\mathbf{W}^{(L)\top}\mathbf{h}^{(L-1)}(\mathbf{x})$. By applying last-layer LA to pre-trained deterministic embedding model, we can convert it to probabilistic embedding model where the last-layer following Gaussian posterior $p(\mathbf{W}^{(L)}\mid\mathcal{D})\approx\mathcal{N}\!\left(\widehat{\mathbf{W}}^{(L)},\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}^{-1}\right)$. The embedding produced by the model becomes a Gaussian random vector obtained by propagating the probabilistic last-layer through the fixed representation:
$\mathbf{z}(\mathbf{x})\sim\mathcal{N}\!\left(\widehat{\mathbf{W}}^{(L)\top}\mathbf{h}^{(L-1)}(\mathbf{x}),\;\mathbf{h}^{(L-1)}(\mathbf{x})^{\top}\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}^{-1}\mathbf{h}^{(L-1)}(\mathbf{x})\right).$
We adopt a diagonal approximation to $\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}$ for computational efficiency, following prior works (Daxberger et al., 2021; Zhdanov et al., 2025). These Gaussian embeddings are then leveraged to construct an uncertainty-aware convolution across multiple embedding models.
3.2 Uncertainty-driven Coefficients for Embedding Convolution
@@
Suppose we have $K$ independent embedding models, each transformed through the procedure in Section 3.1 into a probabilistic embedding model that produces $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$. To integrate these into a single representation, we perform a Gaussian convolution in the embedding space. Formally, we define the convolutional embedding $\mathbf{z}(\mathbf{x})$ as the result of a weighted combination of independent Gaussian variables:
@@
Uncertainty-aware Coefficients
To determine the coefficients $\bm{\pi}_{k}(\mathbf{x})$ for Gaussian convolution, we introduce an uncertainty-aware surrogate loss that captures how reliably the embeddings of positive pairs are expected to align. This formulation enables a principled aggregation of multiple embedding models while explicitly accounting for model-specific epistemic uncertainty.
Text embeddings are often trained with contrastive objectives such as InfoNCE (Oord et al., 2018). For $\ell_{2}$-normalized features, the squared Euclidean distance is directly related to cosine similarity, $\|\mathbf{u}-\mathbf{v}\|^{2}=2(1-\mathbf{u}^{\top}\mathbf{v})$ (Wang and Isola, 2020). These losses also encourage alignment of positive pairs and uniformity on the unit hypersphere. Leveraging this connection, we adopt the squared loss as a principled surrogate for contrastive objectives.
Let $(\mathbf{x},\mathbf{x}^{\prime})$ denote a positive query–document pair. For each model $k$, the corresponding embeddings follow $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x})),\;\mathbf{z}_{k}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x^{\prime}}),\bm{\Sigma}_{k}(\mathbf{x^{\prime}})).$ We define the uncertainty-aware surrogate loss, a multi-task objective that aggregates model-specific embedding errors weighted by uncertainty-aware task coefficients, as:
@@
This surrogate loss decomposes into fidelity and epistemic uncertainty. While $\bm{\mu}_{k}(\mathbf{x})$ and $\bm{\Sigma}_{k}(\mathbf{x})$ are available at retrieval time, computing $\bm{\mu}_{k}(\mathbf{x}^{\prime})$ and $\bm{\Sigma}_{k}(\mathbf{x}^{\prime})$ requires evaluating the embedding model on every document. However, in practical retrieval systems, document embeddings are precomputed once and stored in the index, whereas the query $\mathbf{x}$ arrives only at inference time (Zhou et al., 2022; Zhao et al., 2024). Since the surrogate loss in Eq. 2 contains terms that depend on $\mathbf{x}^{\prime}$, which cannot be recomputed per query, we discard all document-dependent terms and retain only the query-dependent components. This yields a retrieval-feasible approximation of Eq. 2, leading to the following entropy-regularized convex optimization:
$\min_{\bm{\pi}}\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\left(\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\right)-\;T\,\mathrm{H}\!\left(\bm{\pi}(\mathbf{x})\right)\quad\text{s.t.}\;\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=1,\;\bm{\pi}_{k}(\mathbf{x})\geq 0,$ (3)
where $T>0$ controls sensitivity to uncertainty and $\mathrm{H}(\bm{\pi}(\mathbf{x}))$ denotes the Shannon entropy of the coefficient vector $\bm{\pi}(\mathbf{x}):=(\bm{\pi}_{1}(\mathbf{x}),\ldots,\bm{\pi}_{K}(\mathbf{x}))$, encouraging smooth allocations across models and preventing concentrated solutions. From a Bayesian perspective, this entropy regularization is equivalent (up to a constant) to a KL divergence between $\bm{\pi}(\mathbf{x})$ and a uniform prior over models, encoding a preference against overconfident model selection when uncertainty evidence is limited. Such entropy-based regularization has been widely adopted to stabilize soft reweighting and gating mechanisms under heterogeneous or noisy conditions (Sagawa et al., 2019; Nguyen et al., 2024).
The entropy-regularized objective admits a closed-form solution with temperature-controlled coefficients:
$\bm{\pi}_{k}(\mathbf{x};T)\approx\frac{\exp\big(-\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))/T\big)}{\sum_{j=1}^{K}\exp\big(-\operatorname{tr}(\bm{\Sigma}_{j}(\mathbf{x}))/T\big)}.$ (4)
This approximation follows from $\ell_{2}$ normalization, under which $\|\bm{\mu}_{k}(\mathbf{x})\|^{2}$ is approximately constant across models. With this formulation, UEC aggregates multiple embedding models in a query-adaptive and data-dependent manner. Rather than using fixed or global weights, each encoder’s contribution is modulated by its estimated epistemic uncertainty for the query $\mathbf{x}$, assigning lower influence to unreliable models and emphasizing stable representations. Consequently, UEC performs a principled convolution over probabilistic embeddings that adapts to query-wise heterogeneity and distributional shift. Full derivation and discussion are provided in Appendix A.2.